# AI for Just Work: Constructing Diverse Imaginations of AI beyond “Replacing Humans”

### v1

@@ B.1.1 The concept of imagination in social science
Our proposal of focusing on imagination to the path for social and technical changes is influenced by these schools of thoughts in this subsection on the thesis that imagination can shape and guide practice; and our collective agency is reflected in our capacity to reflect and alter our imaginations and make changes in reality accordingly. For example, in Chapter 1 Dead Zones of the Imagination: An Essay on Structural Stupidity of the book The Utopia of Rules: On Technology, Stupidity, and the Secret Joys of Bureaucracy (Graeber, 2015), anthropologist David Graeber describes the nature of imagination and the possibility to “give power to the imagination”:
@@
In his book Envisioning Real Utopias, sociologist Erik Olin Wright proposes the concept of emancipatory social science that “seeks to generate scientific knowledge relevant to the collective project of challenging various forms of human oppression” (Wright, 2010), and provides pathways to ground the imaginations for social change in pragmatic knowledge and understanding about the complexity of social systems and humanity. This work inspires us to base the new imaginations on “diagnosis and critique of the causal processes that generate these harm” (Wright, 2010). The following is an excerpt in Chapter 1 Introduction: Why real utopias? on the relationship between imagination and practice:
@@
In addition to our proposed new imagination, another example of the new imagination of AI is the Indigenous AI project. Named Abundant Intelligences, this new research agenda of AI is based on Indigenous knowledge systems and relational ethics. It reimagines how AI technologies can flourish the Indigenous communities and how AI development can be guided towards a more humane future (Lewis et al., 2024, 2020).
B.1.2 The mainstream imagination of AI
@@ 1. The phenomenon of monoculture in the AI community, and why questioning the objectives of AI matters?
Our inquiry of the imagination of AI echoes with previous rare but important works in the AI community to zoom out and critically inspect the objectives of AI. For example, in his speech Not without us (Weizenbaum, 1986), Joseph Weizenbaum encourages computer scientists and AI experts to critically reflect the purposes of our work:
‘It is a prosaic truth that none of the weapon systems which today threaten murder on a genocidal scale, and whose design, manufacture and sale condemns countless people, especially children, to poverty and starvation, that none of these devices could be developed without the earnest, even enthusiastic, cooperation of computer professionals. It cannot go on without us! Without us the arms race, especially the qualitative arms race, could not advance another step. …… In this context, Artificial Intelligence (AI) comes especially to mind. Many of the technical tasks and problems in this subdiscipline of computer science stimulate the imagination and creativity of technically oriented workers particularly strongly. Goals like making a thinking being out of the computer, giving the computer the ability to understand spoken language, making it possible for the computer to see, goals like these offer nearly irresistible temptations to those among us who have not fully sublimated our playful sandbox fantasies or who mean to satisfy our delusions of omnipotence on the computer stage, i.e., in terms of computer systems. Such tasks are extraordinarily demanding and interesting. Robert Oppenheimer called them sweet. Besides, research projects in these areas are generously funded. The required monies usually come out of the coffers of the military - at least in America. It is enormously tempting and, especially in Artificial Intelligence work, seductively simple to lose or hide oneself in details, in subproblems and their subproblems, and so on. The actual problems on which one works - and which are so generously supported - are disguised and transformed until their representations are more fables, harmless, innocent, lovely fairy tales. …… I don’t quite know whether it is especially computer science or its subdiscipline Artificial Intelligence that has such an enormous affection for euphemism. We speak so spectacularly and so readily of computer systems that understand, that see, decide, make judgments, and so on, without ourselves recognizing our own superficiality and immeasurable naivete with respect to these concepts. And, in the process of so speaking, we anesthetise our ability to evaluate the quality of our work and, what is more important, to identify and become conscious of its end use. …… One can’t escape this state without asking, again and again: “What do I actually do? What is the final application and use of the products of my work?” and ultimately, “am I content or ashamed to have contributed to this use?” ’
Kiri L. Wagstaf in her paper Machine Learning that Matters (Wagstaff, 2012) asks:
@@
In the editorial Machine learning for science and society (Rudin & Wagstaff, 2013), Cynthia Rudin and Kiri L. Wagstaff ask, “what is Machine Learning good for?” And they point out,
@@
Gary Marcus also describes the intensified monoculture in the current AI community, in Chapter 17 Research Into Genuinely Trustworthy AI of his book Taming Silicon Valley: How We Can Ensure That AI Works for Us (Marcus, 2024):
@@
This monoculture tends to exclude diverse works that pose different visions of AI and deviate from the mainstream. For example, Rudin & Wagstaff (2013) mention the issue of monoculture in the publication and career advancement system:
@@
The following excerpt from Jenna Burrell and Jacob Metcalf’s editorial Introduction for the special issue of “Ideologies of AI and the consolidation of power”: Naming power (Burrell & Metcalf, 2024) also reflects the phenomenon that works with different imaginations about the purposes and outcomes of AI could be not welcomed and rejected by mainstream venues in the AI community:
@@
To see why the high-level critical reflection on the objectives and imaginations of AI matters, we would like to refer to a case in obesity science. In her book Soda Science: Making the World Safe for Coca-Cola (Greenhalgh, 2024), anthropologist Susan Greenhalgh tells the story of how industry leader Coca-Cola collaborated with academia to conduct real scientific research that advocated exercise, not calorie restraint, as the priority solution for obesity. This distorted research agenda influenced public health policies on obesity and public understanding on diet and lifestyle in favor of the needs and profits of soda industry. In this case, by exposing and criticizing soda industry’s tactics, it is not meant to deny the value of conducting scientific research on exercise, but to show that setting exercise as the mainstream priority is a tactic to downplay other more important factors that the powerful players don’t encourage us to focus on, which is calorie intake. This is not a solo case. Other studies in the pharmaceutical industry also show how private-oriented power can distort the research agenda, influence the integrity and independence of research, and shape health policies, public understanding on diseases, and medical practice (Sismondo, 2018; Michaels, 2020). The following quotes from the book Ghost-Managed Medicine: Big Pharma’s Invisible Hands (Sismondo, 2018) outline how ordinary research activities can be shaped to server for the private, not public, interests. There are also other notorious examples from the tobacco and energy industries (Brandt, 2009; Oreskes & Conway, 2011).
@@
“Pharmaceutical companies sustain large networks to gather, create, control and disseminate information. They provide the pathways that carry this information, and the energy that makes it move. Through bottlenecks and around curves, knowledge is created, given shape by the channels it navigates. Pharma companies create medical knowledge and move it to where it is most useful; much of it is perfectly ordinary knowledge that happens to support their marketing goals. But because of the companies’ resources, their interests and their levels of control, they become key shapers of almost all medical terrains.” [^9] From Sismondo (2018), Chapter 1 Power and Knowledge in Drug Marketing “Together, the many elements that pharmaceutical companies shape, adjust and assemble constitute markets. These markets are new creations, but because they draw together medical science and health needs they take on an appearance of necessity. They look like entities that have emerged whole from just below the social surface. The goal of pharma’s assemblage marketing is to establish conditions that make specific diagnoses, prescriptions and purchases as obvious and frequent as possible. Ideally, all of the elements of a market can be directed towards the same issues, claims and facts, so that the drugs sell themselves. Pharma companies can then recede into the background, and apply only minimal pressure when needed.” [^10] From Sismondo (2018), Chapter 8 Conclusion: The Haunted Pharmakon
@@
Given the close relationship between AI academia and industry today (Ahmed et al., 2023; Ochigame, 2022; Young et al., 2022; Birhane et al., 2022) and numerous evidence of AI harms (Crawford, 2021; Gray & Suri, 2019; Resnikoff, 2021; Rosenblat, 2018; O’Neil, 2016; Noble, 2018; Costanza-Chock, 2023; Atanasoski & Vora, 2019b; D’Ignazio & Klein, 2020; Chun, 2021; Schaake, 2024; Zuboff, 2019; Eubanks, 2018; Broussard, 2018), these cautionary tales provide significant implications for the AI community: How can we ensure that the current mainstream research agenda in AI is not a remake of the above mistakes in healthcare? While the mainstream research agenda predominantly focuses on productivity, performance, and algorithmic novelty, how can we be certain that we are not trapped by suboptimal priority (Brynjolfsson, 2022), which is analogous to the trap in obesity science to prioritize exercise? How can we be certain that there are not any other more important objectives of AI that the powerful directs us away to exclude from the research agenda, akin to the soda industry’s downplaying of calorie intake? Can we question the validity and ethics of the mainstream research agenda of AI, to reduce the likelihood of being misled by distorted research agenda and being manipulated for unjust purposes?
@@
Although people may tend to assume that technological innovations happen in a free space that allow people to explore different visions and directions at their will, the above phenomenon of monoculture in AI shows that the reality could hardly be the case. This is because technological research and development are not merely personal endeavors carried out by “lone genius inventor” (Matthewman, 2011), they are also social processes and are influenced by complex social factors, such as educational and training systems (Rolnick et al., 2024), funding agencies (Young et al., 2022), career incentives (Rolnick et al., 2024; Rudin & Wagstaff, 2013), and peer review and publication systems (Rolnick et al., 2024; Rudin & Wagstaff, 2013; Andrews et al., 2024; Lipton & Steinhardt, 2019), as pointed out by many previous works in the AI community. These social activities are also embedded in a larger social context where technological innovation cannot be isolated from. Science, Technology, and Society (STS) studies show that “technologies can not be abstracted from the environments which they help to create” (Matthewman, 2011). In this sense, arguing whether technology is value-free, value-neutral, or apolitical is meaningless, because the full lifecycle of technology — its creation, deployment, and use — are always embedded in and inseparable from their social contexts. Therefore, we need to consider the specific social, political, and historical dimensions when understanding and analyzing technology. This is the concept of sociotechnical system in STS that “concerns the ways in which people and technologies are irrevocably entangled across different scales such that human agency is intertwined with technological devices, practices, and infrastructures. …The lens of sociotechnical systems adds the crucial dimension of thinking about how human and technical capacities to act are intertwined. …[It] invites practitioners to interrogate the sociotechnical systems in which one’s project intervenes, which may involve considering the systems upon which the project depends in order to function and also how the output of one’s analysis can instantiate novel sociotechnical arrangements” (Boenig-Liptsin et al., 2022). Failure to consider the social dimension can “render technical interventions ineffective, inaccurate, and sometimes dangerously misguided” (Selbst et al., 2019).
@@
The way technology and society intertwine with each other naturally leads to the second concept that technology and society co-evolve and co-produce each other. The co-production concept requires us to think beyond the technological or social deterministic mindset (Matthewman, 2011). It helps us to understand the source of harms, as well as providing ways for responsible actions in technology and society. As Margarita Boenig-Liptsin, Anissa Tanweer, and Ari Edmundson put it in their paper Data Science Ethos Lifecycle: Interplay of Ethical Thinking and Data Science Practice (Boenig-Liptsin et al., 2022):
@@
Ruha Benjamin also explains the concept of co-production and its relationship with imagination, in her book chapter Introduction: Discriminatory Design, Liberating Imagination (Benjamin, 2019):
@@
Thinking in sociotechnical system and co-production brings the social dimension into the understanding of technology. A prominent idea in sociology is to utilize systemic thinking that regards social system, the whole, larger than a collection of individuals, i.e., a social system contains complex interactions and relationships that cannot be simply reduced to its composing individuals. We will return to this idea in later part on complex systems. As sociologist Allan G. Johnson puts it in his book The Forest and the Trees: Sociology as Life, Practice, and Promise (Johnson, 2014):
@@
“If sociology could teach everyone just one thing with the most profound effect on how we understand social life, it would, I believe, be this: we are always participating in something larger than ourselves, and if we want to understand social life and what happens to people in it, we have to understand what it is that we are participating in and how we are participating in it. In other words, the key to understanding social life is neither just the forest nor just the trees but the forest and the trees and the consequences that result from their dynamic relationship to each other. The larger things we participate in are called social systems, which come in all shapes and sizes. In general, the concept of a ‘system’ refers to any collection of parts or elements that are connected in ways that coalesce into some kind of whole.” [^11] From Johnson (2014) Chapter 1 The Forest, the Trees, and the One Thing
@@
Social systems are like a set of rules in a game that set “paths of least resistance” that affect our thinking and behavior (Johnson, 2014). This systemic perspective allows us to jump out of the individualistic model to understand that the players in the game do not equal to the rules of the game themselves:
@@
“What social life comes down to, then, is a dynamic relationship between social systems and the people who participate in them. Note that people participate in systems without being parts of the systems themselves. … This distinction is easy to lose sight of, but it is crucial. It’s easy to lose sight of because we are so used to thinking solely in terms of individuals. It is crucial because it means that people are not systems and systems are not people, and if we forget that, we are likely to focus on the wrong thing in trying to solve our problems. … Thinking of systems as just people is why members of privileged groups often take it personally when someone points out that their society is racist or sexist. … I cannot make my society or the place where I live suddenly nonracist, but I can decide how to live as a white person in relation to the privileged position of ‘white person’ that I occupy. … I can decide what to do about the consequences that racism produces, whether to be part of the solution or just part of the problem. I do not feel guilty because my country is racist, because the creation of racism in this country was not my doing. But as a white person who participates in that society, I feel responsible to consider what to do about it. The only way to get past the potential for guilt and see how I can make a difference is to realize that the system is not I, nor am I the system. …… For its part, a system affects how we think, feel, and behave as participants. It does this not only through the general process of socialization but also by laying out paths of least resistance in social situations. At any given moment, we could do an almost infinite number of things, but we typically do not realize this and see only a narrow range of possibilities. What the range looks like depends on the system we are in. ……. This is why people might laugh at racist or heterosexist jokes even when they make them feel uncomfortable—in that situation, to not laugh and risk being ostracized by everyone may make them feel even more uncomfortable. The easiest—although not necessarily easy—choice is to go along. This does not mean we have to go along or that we will, only that if we do go along, we’ll run into less resistance than if we don’t. … When we can identify how a system is organized, we can see what is likely to result if people follow the paths of least resistance. We know, for example, where the game of Monopoly is going just by reading the rules of the game. We don’t have to know anything about the individuals who play it, except the likelihood that most of them will follow the path of least resistance most of the time. … People are not systems and systems are not people, which means that social life can produce horrible or wonderful consequences without necessarily meaning that the people who participate are horrible or wonderful themselves. Good people participate in systems that produce bad consequences all the time. …… Getting clear about the relationship between individuals and social systems can dramatically alter how we see potentially painful issues and ourselves in relation to them. This is especially true for people in privileged groups who otherwise resist looking at the nature and consequences of privilege. Their defensive resistance is probably the biggest single barrier to ending privilege and oppression. Most of the time resistance happens because, like everyone else, people in privileged groups are stuck in an individualistic model of the world and cannot see how to acknowledge white privilege as a fact of social life without also feeling personally to blame for it. And the people who are most likely to feel this way are often the ones who are otherwise most open to doing something to make things better. ” [^12] From Johnson (2014) Chapter 1 The Forest, the Trees, and the One Thing
@@
Another important concept is social structure, which is “[t]he patterns of relationships and distributions that characterize the organization of a social system. Relationships connect various elements of a system (such as social statuses) to one another and to the system itself. Distributions include valued resources and rewards, such as power and income, and the distribution of people among social statuses. Structure can also refer to relationships and distributions among systems” (Johnson, 2014). An important asymmetric relationship in the social structure lies in the concept of power, which is defined as “the asymmetric capacity to structure or alter the behavior of others” (Boenig-Liptsin et al., 2022).
@@
“Having a set of cultural beliefs allows us to live with a taken-for-granted sense of how things are and to treat the ‘facts’ of our existence as obvious. What we call ‘obvious,’ however, is not necessarily what is true. It is only assumed to be true beyond doubt in a particular culture. Without a sense of the obvious, social life loses its predictability, and we lose our basis for feeling secure, but the obvious also blinds us to the possibility that what is ‘obviously’ true may be false. …… I can expand my freedom only by liberating myself from the narrow range of choices that my culture—that any culture—offers the people who participate in it. To do this, I need to step outside the cultural framework I am used to so that I can see it as a framework, as one possibility among many. Stepping outside is an important part of what sociological practice is about, and such concepts as culture, beliefs, and values are important tools used in the process, for they point to what we are stepping outside of.” [^13] From Johnson (2014) Chapter 2 Culture: Symbols, Ideas, and the Stuff of Life “Sociology … can change how we see and experience reality by revealing assumptions and understandings that underlie everyday life. Many of these are rarely spoken or otherwise made explicit, and yet they operate in powerful ways to shape our perceptions, thoughts, feelings, and behavior. By going beneath the surface, sociology reveals a deeper and more complex reality.” [^14] From Johnson (2014) Chapter 6 Things Are Not What They Seem

### v2

@@ D.1.1 The concept of imagination in social science
Our proposal of focusing on imagination to the path for social and technical changes is influenced by these schools of thoughts in this subsection on the thesis that imagination can shape and guide practice; and our collective agency is reflected in our capacity to reflect and alter our imaginations and make changes in reality accordingly. For example, in Chapter 1 Dead Zones of the Imagination: An Essay on Structural Stupidity of the book The Utopia of Rules: On Technology, Stupidity, and the Secret Joys of Bureaucracy [55], anthropologist David Graeber describes the nature of imagination and the possibility to “give power to the imagination”:
@@
In his book Envisioning Real Utopias, sociologist Erik Olin Wright proposes the concept of emancipatory social science that “seeks to generate scientific knowledge relevant to the collective project of challenging various forms of human oppression” [109], and provides pathways to ground the imaginations for social change in pragmatic knowledge and understanding about the complexity of social systems and humanity. This work inspires us to base the new imaginations on “diagnosis and critique of the causal processes that generate these harm” [109]. The following is an excerpt in Chapter 1 Introduction: Why real utopias? on the relationship between imagination and practice:
@@
In addition to our proposed new imagination, another example of the new imagination of AI is the Indigenous AI project. Named Abundant Intelligences, this new research agenda of AI is based on Indigenous knowledge systems and relational ethics. It reimagines how AI technologies can flourish the Indigenous communities and how AI development can be guided towards a more humane future [70, 69].
D.1.2 The mainstream imagination of AI
@@ 1. The phenomenon of monoculture in the AI community, and why questioning the objectives of AI matters?
Our inquiry of the imagination of AI echoes with previous rare but important works in the AI community to zoom out and critically inspect the objectives of AI. For example, in his speech Not without us [106], Joseph Weizenbaum encourages computer scientists and AI experts to critically reflect the purposes of our work:
‘It is a prosaic truth that none of the weapon systems which today threaten murder on a genocidal scale, and whose design, manufacture and sale condemns countless people, especially children, to poverty and starvation, that none of these devices could be developed without the earnest, even enthusiastic, cooperation of computer professionals. It cannot go on without us! Without us the arms race, especially the qualitative arms race, could not advance another step. …… In this context, Artificial Intelligence (AI) comes especially to mind. Many of the technical tasks and problems in this subdiscipline of computer science stimulate the imagination and creativity of technically oriented workers particularly strongly. Goals like making a thinking being out of the computer, giving the computer the ability to understand spoken language, making it possible for the computer to see, goals like these offer nearly irresistible temptations to those among us who have not fully sublimated our playful sandbox fantasies or who mean to satisfy our delusions of omnipotence on the computer stage, i.e., in terms of computer systems. Such tasks are extraordinarily demanding and interesting. Robert Oppenheimer called them sweet. Besides, research projects in these areas are generously funded. The required monies usually come out of the coffers of the military - at least in America. It is enormously tempting and, especially in Artificial Intelligence work, seductively simple to lose or hide oneself in details, in subproblems and their subproblems, and so on. The actual problems on which one works - and which are so generously supported - are disguised and transformed until their representations are more fables, harmless, innocent, lovely fairy tales. …… I don’t quite know whether it is especially computer science or its subdiscipline Artificial Intelligence that has such an enormous affection for euphemism. We speak so spectacularly and so readily of computer systems that understand, that see, decide, make judgments, and so on, without ourselves recognizing our own superficiality and immeasurable naivete with respect to these concepts. And, in the process of so speaking, we anesthetise our ability to evaluate the quality of our work and, what is more important, to identify and become conscious of its end use. …… One can’t escape this state without asking, again and again: “What do I actually do? What is the final application and use of the products of my work?” and ultimately, “am I content or ashamed to have contributed to this use?" ’
Kiri L. Wagstaf in her paper Machine Learning that Matters [104] asks:
@@
In the editorial Machine learning for science and society [97], Cynthia Rudin and Kiri L. Wagstaff ask, “what is Machine Learning good for?” And they point out,
@@
Gary Marcus also describes the intensified monoculture in the current AI community, in Chapter 17 Research Into Genuinely Trustworthy AI of his book Taming Silicon Valley: How We Can Ensure That AI Works for Us [73]:
@@
This monoculture tends to exclude diverse works that pose different visions of AI and deviate from the mainstream. For example, Rudin and Wagstaff, [97] mention the issue of monoculture in the publication and career advancement system:
@@
The following excerpt from Jenna Burrell and Jacob Metcalf’s editorial Introduction for the special issue of “Ideologies of AI and the consolidation of power”: Naming power [31] also reflects the phenomenon that works with different imaginations about the purposes and outcomes of AI could be not welcomed and rejected by mainstream venues in the AI community:
@@
To see why the high-level critical reflection on the objectives and imaginations of AI matters, we would like to refer to a case in obesity science. In her book Soda Science: Making the World Safe for Coca-Cola [57], anthropologist Susan Greenhalgh tells the story of how industry leader Coca-Cola collaborated with academia to conduct real scientific research that advocated exercise, not calorie restraint, as the priority solution for obesity. This distorted research agenda influenced public health policies on obesity and public understanding on diet and lifestyle in favor of the needs and profits of soda industry. In this case, by exposing and criticizing soda industry’s tactics, it is not meant to deny the value of conducting scientific research on exercise, but to show that setting exercise as the mainstream priority is a tactic to downplay other more important factors that the powerful players don’t encourage us to focus on, which is calorie intake. This is not a solo case. Other studies in the pharmaceutical industry also show how private-oriented power can distort the research agenda, influence the integrity and independence of research, and shape health policies, public understanding on diseases, and medical practice [100, 76]. The following quotes from the book Ghost-Managed Medicine: Big Pharma’s Invisible Hands [100] outline how ordinary research activities can be shaped to server for the private, not public, interests. There are also other notorious examples from the tobacco and energy industries [26, 87].
@@
“Pharmaceutical companies sustain large networks to gather, create, control and disseminate information. They provide the pathways that carry this information, and the energy that makes it move. Through bottlenecks and around curves, knowledge is created, given shape by the channels it navigates. Pharma companies create medical knowledge and move it to where it is most useful; much of it is perfectly ordinary knowledge that happens to support their marketing goals. But because of the companies’ resources, their interests and their levels of control, they become key shapers of almost all medical terrains.” [^11] From Sismondo, [100], Chapter 1 Power and Knowledge in Drug Marketing “Together, the many elements that pharmaceutical companies shape, adjust and assemble constitute markets. These markets are new creations, but because they draw together medical science and health needs they take on an appearance of necessity. They look like entities that have emerged whole from just below the social surface. The goal of pharma’s assemblage marketing is to establish conditions that make specific diagnoses, prescriptions and purchases as obvious and frequent as possible. Ideally, all of the elements of a market can be directed towards the same issues, claims and facts, so that the drugs sell themselves. Pharma companies can then recede into the background, and apply only minimal pressure when needed.” [^12] From Sismondo, [100], Chapter 8 Conclusion: The Haunted Pharmakon
@@
Given the close relationship between AI academia and industry today [10, 85, 111, 20] and numerous evidence of AI harms [38, 56, 93, 96, 86, 84, 37, 15, 41, 35, 98, 114, 43, 27], these cautionary tales provide significant implications for the AI community: How can we ensure that the current mainstream research agenda in AI is not a remake of the above mistakes in healthcare? While the mainstream research agenda predominantly focuses on productivity, performance, and algorithmic novelty, how can we be certain that we are not trapped by suboptimal priority [28], which is analogous to the trap in obesity science to prioritize exercise? How can we be certain that there are not any other more important objectives of AI that the powerful directs us away to exclude from the research agenda, akin to the soda industry’s downplaying of calorie intake? Can we question the validity and ethics of the mainstream research agenda of AI, to reduce the likelihood of being misled by distorted research agenda and being manipulated for unjust purposes?
@@
Although people may tend to assume that technological innovations happen in a free space that allow people to explore different visions and directions at their will, the above phenomenon of monoculture in AI shows that the reality could hardly be the case. This is because technological research and development are not merely personal endeavors carried out by “lone genius inventor” [74], they are also social processes and are influenced by complex social factors, such as educational and training systems [95], funding agencies [111], career incentives [95, 97], and peer review and publication systems [95, 97, 13, 71], as pointed out by many previous works in the AI community. These social activities are also embedded in a larger social context where technological innovation cannot be isolated from. Science, Technology, and Society (STS) studies show that “technologies can not be abstracted from the environments which they help to create” [74]. In this sense, arguing whether technology is value-free, value-neutral, or apolitical is meaningless, because the full lifecycle of technology — its creation, deployment, and use — are always embedded in and inseparable from their social contexts. Therefore, we need to consider the specific social, political, and historical dimensions when understanding and analyzing technology. This is the concept of sociotechnical system in STS that “concerns the ways in which people and technologies are irrevocably entangled across different scales such that human agency is intertwined with technological devices, practices, and infrastructures. …The lens of sociotechnical systems adds the crucial dimension of thinking about how human and technical capacities to act are intertwined. …[It] invites practitioners to interrogate the sociotechnical systems in which one’s project intervenes, which may involve considering the systems upon which the project depends in order to function and also how the output of one’s analysis can instantiate novel sociotechnical arrangements” [24]. Failure to consider the social dimension can “render technical interventions ineffective, inaccurate, and sometimes dangerously misguided” [99].
@@
The way technology and society intertwine with each other naturally leads to the second concept that technology and society co-evolve and co-produce each other. The co-production concept requires us to think beyond the technological or social deterministic mindset [74]. It helps us to understand the source of harms, as well as providing ways for responsible actions in technology and society. As Margarita Boenig-Liptsin, Anissa Tanweer, and Ari Edmundson put it in their paper Data Science Ethos Lifecycle: Interplay of Ethical Thinking and Data Science Practice [24]:
@@
Ruha Benjamin also explains the concept of co-production and its relationship with imagination, in her book chapter Introduction: Discriminatory Design, Liberating Imagination [17]:
@@
Thinking in sociotechnical system and co-production brings the social dimension into the understanding of technology. A prominent idea in sociology is to utilize systemic thinking that regards social system, the whole, larger than a collection of individuals, i.e., a social system contains complex interactions and relationships that cannot be simply reduced to its composing individuals. We will return to this idea in later part on complex systems. As sociologist Allan G. Johnson puts it in his book The Forest and the Trees: Sociology as Life, Practice, and Promise [63]:
@@
“If sociology could teach everyone just one thing with the most profound effect on how we understand social life, it would, I believe, be this: we are always participating in something larger than ourselves, and if we want to understand social life and what happens to people in it, we have to understand what it is that we are participating in and how we are participating in it. In other words, the key to understanding social life is neither just the forest nor just the trees but the forest and the trees and the consequences that result from their dynamic relationship to each other. The larger things we participate in are called social systems, which come in all shapes and sizes. In general, the concept of a ‘system’ refers to any collection of parts or elements that are connected in ways that coalesce into some kind of whole.” [^13] From Johnson, [63] Chapter 1 The Forest, the Trees, and the One Thing
@@
Social systems are like a set of rules in a game that set “paths of least resistance” that affect our thinking and behavior [63]. This systemic perspective allows us to jump out of the individualistic model to understand that the players in the game do not equal to the rules of the game themselves:
@@
“What social life comes down to, then, is a dynamic relationship between social systems and the people who participate in them. Note that people participate in systems without being parts of the systems themselves. … This distinction is easy to lose sight of, but it is crucial. It’s easy to lose sight of because we are so used to thinking solely in terms of individuals. It is crucial because it means that people are not systems and systems are not people, and if we forget that, we are likely to focus on the wrong thing in trying to solve our problems. … Thinking of systems as just people is why members of privileged groups often take it personally when someone points out that their society is racist or sexist. … I cannot make my society or the place where I live suddenly nonracist, but I can decide how to live as a white person in relation to the privileged position of ‘white person’ that I occupy. … I can decide what to do about the consequences that racism produces, whether to be part of the solution or just part of the problem. I do not feel guilty because my country is racist, because the creation of racism in this country was not my doing. But as a white person who participates in that society, I feel responsible to consider what to do about it. The only way to get past the potential for guilt and see how I can make a difference is to realize that the system is not I, nor am I the system. …… For its part, a system affects how we think, feel, and behave as participants. It does this not only through the general process of socialization but also by laying out paths of least resistance in social situations. At any given moment, we could do an almost infinite number of things, but we typically do not realize this and see only a narrow range of possibilities. What the range looks like depends on the system we are in. ……. This is why people might laugh at racist or heterosexist jokes even when they make them feel uncomfortable—in that situation, to not laugh and risk being ostracized by everyone may make them feel even more uncomfortable. The easiest—although not necessarily easy—choice is to go along. This does not mean we have to go along or that we will, only that if we do go along, we’ll run into less resistance than if we don’t. … When we can identify how a system is organized, we can see what is likely to result if people follow the paths of least resistance. We know, for example, where the game of Monopoly is going just by reading the rules of the game. We don’t have to know anything about the individuals who play it, except the likelihood that most of them will follow the path of least resistance most of the time. … People are not systems and systems are not people, which means that social life can produce horrible or wonderful consequences without necessarily meaning that the people who participate are horrible or wonderful themselves. Good people participate in systems that produce bad consequences all the time. …… Getting clear about the relationship between individuals and social systems can dramatically alter how we see potentially painful issues and ourselves in relation to them. This is especially true for people in privileged groups who otherwise resist looking at the nature and consequences of privilege. Their defensive resistance is probably the biggest single barrier to ending privilege and oppression. Most of the time resistance happens because, like everyone else, people in privileged groups are stuck in an individualistic model of the world and cannot see how to acknowledge white privilege as a fact of social life without also feeling personally to blame for it. And the people who are most likely to feel this way are often the ones who are otherwise most open to doing something to make things better. ” [^14] From Johnson, [63] Chapter 1 The Forest, the Trees, and the One Thing
@@
Another important concept is social structure, which is “[t]he patterns of relationships and distributions that characterize the organization of a social system. Relationships connect various elements of a system (such as social statuses) to one another and to the system itself. Distributions include valued resources and rewards, such as power and income, and the distribution of people among social statuses. Structure can also refer to relationships and distributions among systems” [63]. An important asymmetric relationship in the social structure lies in the concept of power, which is defined as “the asymmetric capacity to structure or alter the behavior of others” [24].
@@
“Having a set of cultural beliefs allows us to live with a taken-for-granted sense of how things are and to treat the ‘facts’ of our existence as obvious. What we call ‘obvious,’ however, is not necessarily what is true. It is only assumed to be true beyond doubt in a particular culture. Without a sense of the obvious, social life loses its predictability, and we lose our basis for feeling secure, but the obvious also blinds us to the possibility that what is ‘obviously’ true may be false. …… I can expand my freedom only by liberating myself from the narrow range of choices that my culture—that any culture—offers the people who participate in it. To do this, I need to step outside the cultural framework I am used to so that I can see it as a framework, as one possibility among many. Stepping outside is an important part of what sociological practice is about, and such concepts as culture, beliefs, and values are important tools used in the process, for they point to what we are stepping outside of.” [^15] From Johnson, [63] Chapter 2 Culture: Symbols, Ideas, and the Stuff of Life “Sociology … can change how we see and experience reality by revealing assumptions and understandings that underlie everyday life. Many of these are rarely spoken or otherwise made explicit, and yet they operate in powerful ways to shape our perceptions, thoughts, feelings, and behavior. By going beneath the surface, sociology reveals a deeper and more complex reality.” [^16] From Johnson, [63] Chapter 6 Things Are Not What They Seem